OpenStudy (anonymous):

the length of a rectangle is 6 time its width. The perimeter is 98 ft. What is its length and width?

OpenStudy (anonymous):

98/2 = 49 [1 length + 1 width] l = 6w , so 49 / (6 + 1) = 7 which is the length and 7*6 = 42

OpenStudy (anonymous):